Polynomial Inequalities:

4) x^2 + 8x > 0

show your number line test and state answer in interval notation


1
Expert's answer
2021-09-20T15:59:40-0400

We may rewrite the inequality in form

"x(x+8) > 0."

Next, we solve the equation

"x(x+8) = 0."

The roots are equal to 0 and -8. We place these roots on the line and mark the sign of the left hand of the inequality.


At "(-\infty, -8)" "x<0, x+8< 0 \; \Rightarrow x(x+8) >0,"

at "(-8,0)\;\; \;\;x<0, x+8> 0 \; \Rightarrow x(x+8) <0,"

at "(0,+\infty)\;\; \;\;x>0, x+8> 0 \; \Rightarrow x(x+8) >0,"





The answer will be "x\in(-\infty, -8) \cup (0,+\infty)"